# Dataset Distillation-based Hybrid Federated Learning on Non-IID Data

### v2

@@ III-B Federated Learning
Consider a classic FL scenario with $N$ participants collaboratively training a global model. Each participant possesses a local dataset, represented as $\mathcal{D}_{i}=\left\{\left(X_{i},y_{i}\right)\right\},i=1,\cdots,N$, where $X_{i}\in\mathbb{R}^{n_{i}\times d}$ and $y_{i}\in\mathbb{R}^{n_{i}\times n_{c}}$ denote the tuple of input data points and their corresponding labels. $n=\sum_{i=1}^{N}{n_{i}}$, where $n_{i}=|\mathcal{D}_{i}|$ is the number of samples in $\mathcal{D}_{i}$. The objective of collaborative learning is to solve the following distributed parallel optimization problem
[…]
@@
In the process of training, each participating client trains the local model based on their local dataset and then transmits the local models to a server for global model aggregation. Specifically, in round $t$, the client $i$ trains its local model $\omega_{i}^{t}$ by minimizing the local loss function $f_{i}\left(\omega\right)$ using the gradient descent method. Subsequently, the server updates the global model $\omega^{t}$ through aggregation rules, e.g., FedAvg. If client data are IID, the above method can provide the optimal solution of (2). while if the client data are non-IID, the obtained global model may be drifted from the optimum. The communication cost of the above process comes from the local model upload and the global model download. By referring to Ref. [29], the communication cost of the widely used FedAvg is
$Cost_{fedavg}\,\,=\,\,N\cdot\left(2T-1\right)\cdot|\omega|\cdot B_{1},$ (3)
@@
TABLE I: Important Notations
@@
$\tilde{\mathcal{D}_{i}}$ distilled dataset on client $i$
@@
$\mathcal{D}_{h}$ whole dataset on the $h$-th cluster head
$S_{i}$ soft label matrix generated by client $i$
@@
$E_{1}$ number of local update steps in pre-train at label knowledge collection stage
$E_{2}$ number of local update steps in local training at model training stage
@@
Figure 1: Overall pipeline of HFLDD. Different colors represent different label classes. There are $N$ clients divided into $H$ heterogeneous clusters, where each cluster’s clients have distinct label classes.
@@ III-C Problem Statement
In this paper, we consider a typical scenario of distributed learning with non-IID data, i.e., label distribution skew [46]. Our focus lies on addressing the issue of quantity-based label imbalance, where each client owns data with a fixed number of labels. Supposing $N_{c}$ denotes the total number of label classes, each client owns data with $C$ different labels. Data labels among different clients could overlap. If $C=N_{c}$, the data labels among all clients are the same, which can be regarded as quantity-based label balance. If $C<N_{c}$, the data labels among all clients are imbalanced. If $C=1$, it represents the worst-case scenario of quantity-based label imbalance, where the local dataset on each client contains only a single label. The quantity-based label imbalance has been verified to greatly decrease the performance of model training [14, 47].
Instead of traditional FL, we will design a new HFL framework that integrates dataset distillation to create approximately label-balanced local datasets, thereby alleviating global model drift caused by non-IID data. Furthermore, it should be noted that clients can communicate with each other.
@@ IV The Proposed HFLDD
In this section, we will first introduce the key idea of the proposed framework HFLDD. Then, we will introduce the details of each part and finally provide an analysis of the communication cost of HFLDD.
@@ IV-A Overview
    Label Knowledge Collection (Step 1): Each client independently trains a local model using its private non-IID dataset. These locally trained models are then used to infer soft labels on a globally available auxiliary dataset. The resulting soft-label matrices are uploaded to the server. This stage allows the server to gather label distribution information without accessing the raw private data, thus preserving client privacy.
@@
    IID Dataset Generation (Step 4): Within each heterogeneous cluster, all clients except the cluster head perform dataset distillation to generate a compact synthetic dataset that approximates their local data distribution. These distilled datasets are sent to the cluster head. The head aggregates all received distilled data, including its own data, to form a more balanced, approximately IID dataset. This step helps mitigate statistical heterogeneity before training begins.
@@
    Model Training (Step 5): The server distributes the current global model to all cluster heads. Each head updates the model using the locally aggregated IID-like dataset and sends the updated model back to the server. The server aggregates the updated models using standard FL techniques such as FedAvg. This process is repeated for multiple rounds until model convergence.
Figure 2: An overview of heterogeneous client clustering. Each circle represents a client, with different colors representing different data labels. Initially, the server collects the soft labels uploaded by clients and calculates the similarity matrix. Then the server groups the clients into homogeneous clusters through K-Means, and reorganizes them into heterogeneous clusters through random sampling.
@@ IV-B Details
The details of the implementation of each part in HFLDD are introduced in this subsection. Table I summarizes the important notations.
@@ IV-B1 Label knowledge collection
Each client first trains a local model on their own dataset $\mathcal{D}_{i}$. Since the trained local model is driven by the client dataset, it learns relevant knowledge about the local data. Subsequently, we introduce a finite and globally available dataset $\mathcal{D}_{g}=\{\left(X_{g},y_{g}\right)\}$  [^1] The global dataset $\mathcal{D}_{g}$ used for soft label generation can be a small, publicly available dataset related to the same task but it is not necessarily representative of the client data, where $X_{g}\in\mathbb{R}^{|\mathcal{D}_{g}|\times d}$, $y_{g}\in\mathbb{R}^{|\mathcal{D}_{g}|\times N_{c}}$ and $|\mathcal{D}_{g}|$ is the number of samples in $\mathcal{D}_{g}$. The global dataset has a distribution different from the datasets of the clients and is used to generate soft labels for each client. Specifically, each client uses a globally available dataset as input for its locally pre-trained model to generate a two-dimensional tensor $S_{i}\in\mathbb{R}^{|\mathcal{D}_{g}|\times N_{c}},\forall i=1,\cdots,N$. This tensor represents the soft labels of the global dataset based on the local model of the client $i$. ${S}_{i,k}^{c}$ denotes the probability that the sample $k$ belongs to class $c$. It has been verified that the above soft labels can reflect the knowledge of the label of the local dataset [32, 48]. Therefore, after receiving the soft labels from the clients, the server obtains the label knowledge of each client without knowing the original dataset, which prevents data privacy leakage.
@@ IV-B2 Heterogeneous Client Clustering
Based on the similarity matrix, the server employs the K-Means algorithm to cluster all the clients, by treating each row of the similarity matrix as a data point. This process groups clients with similar knowledge of labels into homogeneous groups $\mathcal{H}_{o}=\left\{\mathcal{H}_{o,1},\ldots,\mathcal{H}_{o,K}\right\}$. However, our goal is to construct heterogeneous clusters $\mathcal{H}_{e}=\left\{\mathcal{H}_{e,1},\ldots,\mathcal{H}_{e,H}\right\}$ with approximately intracluster IID data. Therefore, to construct heterogeneous clusters, we randomly select one client from each of the homogeneous clusters and form a new cluster with them. The details are shown in Algorithm 1. Through this sampling process, we ensure that the newly formed clusters contain clients with different label knowledge. Fig. 2 illustrates the overall process of heterogeneous client clustering. Each client’s local data contains only one label, which makes the data among different clients severe non-IID. After clustering, each homogeneous cluster contains only one label, while each heterogeneous cluster contains labels from all clients.
@@ IV-B3 IID dataset generation
In each heterogeneous cluster, the cluster head acts as an agent for all members of the cluster, conducts local model training, and communicates with the server. Each cluster member, except the head, sends a distilled data of much smaller size to the cluster head. In this work, clients employ the KIP method to obtain distilled data, as shown in Section III-A. For the $h$-th heterogeneous cluster, when the cluster head aggregates the distilled data from the entire cluster, it obtains a dataset $\mathcal{D}_{h}$, whose distribution approximates the global distribution. We transform the problem into a learning process between the server and a number of cluster heads with approximately IID data.
@@ IV-B4 Model Training
As we mentioned above, after transforming our problem into distributed learning with IID data, we can apply the classical FL method, such as FedAvg[11], for model training. In the communication round $t$, the server broadcasts the global model parameters $\omega^{t}$ to all heads of the connected cluster. Subsequently, the cluster head $h$ updates its local parameters to $\omega_{h}^{t}\,\,=\,\,\omega^{t}$, and performs $E_{2}$ local updates in $\mathcal{D}_{h}$. At the $(k+1)$-th step of local update, the local model parameters of cluster head $h$ are updated as
$\omega_{h}^{t,k+1}=\omega_{h}^{t,k}-\eta\nabla F_{h}\left(\omega_{h}^{t,k},\xi_{h}^{t,k}\right),$ (5)
where $h=1,\cdots,H$, $k=0,\cdots,E_{2}-1$, $\xi_{h}^{t,k}$ denotes a randomly selected batch from $\mathcal{D}_{h}$, $\eta$ denotes the learning rate and $\nabla F_{h}\left(\cdot\right)$ denotes the gradient of the local loss function. Finally, the server aggregates the updated local models $\omega_{1}^{t,E_{2}},...,\omega_{H}^{t,E_{2}}$ from all the heads of the cluster to generate a new global model.
The general process of the proposed HFLDD is illustrated in Algorithm 2.
@@
10:  for client $\text{client }i=1$ to $N$ do
11:     for client $\text{client }j=1$ to $N$ do
@@
15:  $\mathcal{H}_{o}=\left\{\mathcal{H}_{o,1},\ldots,\mathcal{H}_{o,K}\right\}\,\,\longleftarrow\,\,\text{K-Means}(K,M)$
16:  $\mathcal{H}_{e}\,\,\longleftarrow\,\,\mathbf{ClusterSampling}(K,\mathcal{H}_{o})$
17:  $\mathcal{A}=\emptyset$
@@ V-A Convergence Analysis
In HFLDD, the global training process is organized around a set of cluster heads ${h=1,\dots,H}$, each of which holds a hybrid dataset consisting of its own local data and the distilled data received from its cluster members. Each cluster head independently performs model training using this hybrid dataset, which approximates the distribution of the entire cluster. The convergence behavior of HFLDD can be analyzed by extending the standard convergence theory of FedAvg with non-IID data. HFLDD maintains the following foundational assumptions commonly adopted in FL:
@@
$||\nabla f_{h}(\omega)-\nabla f_{h}(\upsilon)||^{2}\leq L||\omega-\upsilon||^{2},\forall h=1,\dots,H.$ (6)
@@
$f_{h}(\upsilon)\geq f_{h}(\omega)+(\upsilon-\omega)^{T}\nabla f_{h}(\omega)+\frac{\mu}{2}||\upsilon-\omega||^{2},\forall h=1,\dots,H.$ (7)
@@
$\mathbb{E}_{\xi_{h}\sim\mathcal{D}_{h}}\left[||\nabla f_{h}(w,\xi_{h})-\nabla f_{h}(w)||^{2}\right]\leq\sigma_{h}^{2}$ (9)
@@
$\mathbb{E}_{\xi_{h}\sim\mathcal{D}_{h}}\left[||\nabla f_{h}(w,\xi_{h})||^{2}\right]\leq G^{2},~{}\forall h=1,\dots,H.$ (10)
@@
Assumption 6 (Bounded distillation error): The approximation error between the empirical loss on real dataset and the distilled data at each cluster head is bounded by:
@@
Under Assumptions 1–6, if the learning rate follows a diminishing schedule $\eta_{t}=\frac{2}{\mu(\tau+t)}$, with $\tau=\max\left\{8\frac{L}{\mu},\ E_{2}\right\}-1$, and all cluster heads participate in each communication round, then after $t$ rounds of communication, the global model $\omega_{T}$ satisfies:
$\mathbb{E}[f(\omega^{t})]-f^{*}$ $\leq\frac{L}{t+\tau}\left(\frac{2Q}{\mu^{2}}+\frac{\tau+1}{2}||\omega^{0}-\omega^{*}||^{2}\right),$ (13)
@@
In HFLDD, only the cluster heads participant in global model training. It can be regarded as standard FedAvg on a cluster-level hybrid dataset $\mathcal{D}_{h}$. Based on the convergence analysis of FedAvg under non-IID conditions, as shown in [12], the convergence bound of HFLDD satisfies:
$\mathbb{E}[f(\omega^{t})]-f^{*}$ $\leq\frac{L}{t+\tau}\left(\frac{2B}{\mu^{2}}+\frac{\tau+1}{2}||\omega^{0}-\omega^{*}||^{2}\right),$ (15)
@@ V-B Communication Cost
Now, we analyze the total communication cost of HFLDD. Considering that the initialization of the server model can be sent as a random seed [29], we can ignore the communication cost of sending random seeds. Thus, we only consider the resources consumed for uploading soft labels to the server at the stage of label knowledge collection, transmitting the distilled data within each cluster at the stage of IID dataset generation, and communication between the server and cluster heads at the stage of model training.
Let $a_{h}$ denote the head of the heterogeneous cluster $h$, and $\left|\tilde{\mathcal{D}_{i}}\right|$ denote the size of the distilled dataset on the client $i$. Suppose that transmitting each parameter incurs a communication overhead of $B_{1}$ (usually 32 bits), and transmitting each distilled data incurs a communication overhead of $B_{2}$ (1 × 8 bits or 3 × 8 bits for each grayscale or RGB pixel). The bit length $B_{2}$ reflects the transmission cost of each image sample, and no additional compression beyond the inherent reduction via dataset distillation is applied. For $N$ clients, the communication cost for uploading the soft label is $N\cdot|\mathcal{D}_{g}|\cdot N_{c}\cdot B_{1}$. The total communication cost to transmit the distilled data within each cluster is $\sum_{h=1}^{H}{\sum_{\begin{subarray}{c}i\in\mathcal{H}_{e,h}\\ i\neq a_{h}\end{subarray}}{\left|\tilde{\mathcal{D}_{i}}\right|}}\cdot B_{2}$. In the model training stage, the total cost of model communication between the server and the heads of the cluster in $T$ rounds is $H\cdot|\omega|\cdot(2T-1)\cdot B_{1}$. Consequently, the total communication cost of HFLDD is
$Cost_{hfldd}$ $=N\cdot|\mathcal{D}_{g}|\cdot N_{c}\cdot B_{1}$
$~{}~{}~{}+\sum_{h=1}^{H}{\sum_{\begin{subarray}{c}i\in\mathcal{H}_{e,h}\\ i\neq a_{h}\end{subarray}}{\left|\tilde{\mathcal{D}_{i}}\right|}}\cdot B_{2}$

### v3

@@ III-B Federated Learning
Consider a classic FL scenario with $N$ participants collaboratively training a global model. Each participant possesses a local dataset, represented as $\mathcal{D}_{i}=\left\{\left(X_{i},\mathrm{y}_{i}\right)\right\},i=1,\cdots,N$, where $X_{i}\in\mathbb{R}^{n_{i}\times d}$ and $\mathrm{y}_{i}\in\mathbb{R}^{n_{i}\times n_{c}}$ denote the tuple of input data points and their corresponding labels; $n=\sum_{i=1}^{N}{n_{i}}$, where $n_{i}=|\mathcal{D}_{i}|$ is the number of samples in $\mathcal{D}_{i}$. The objective of collaborative learning is to solve the following distributed parallel optimization problem
$\underset{\omega}{\min}f\left(\omega\right)\triangleq\underset{\omega}{\min}\sum_{i=1}^{N}{\frac{n_{i}}{n}f_{i}\left(\omega\right)},$ (2)
@@
In the process of training, each participating client trains the local model based on their local dataset and then transmits the local models to a server for global model aggregation. Specifically, in round $t$, client $i$ trains its local model $\omega_{i}^{t}$ by minimizing the local loss function $f_{i}\left(\omega\right)$ using the gradient descent method. Subsequently, the server updates the global model $\omega^{t}$ through aggregation rules, e.g., FedAvg. If client data are IID, the above method can provide the optimal solution of (2). If the client data are non-IID, the obtained global model may be drifted from the optimum. The communication cost of the above process comes from the local model upload and the global model download. By referring to [zhou2020distilled], the communication cost of the widely used FedAvg is
$\mathcal{C}_{\text{FedAvg}}\,\,=\,\,N\cdot\left(2T-1\right)\cdot|\omega|\cdot B_{1},$ (3)
@@
TABLE II: Important Notations
@@
$\tilde{\mathcal{D}_{i}}$ distilled data on client $i$
@@
$\mathcal{D}_{h}$ hybrid dataset on the $h$-th cluster head
$S_{i}$ soft labels generated by client $i$
@@
$E_{1}$ number of epochs in pre-train at label knowledge collection stage
$E_{2}$ number of epochs in local training at model training stage
@@
Figure 1: Overall pipeline of HFLDD. Different colors represent different label classes. Each client first performs local training and uploads soft labels to the server. The server conducts heterogeneous client clustering based on the similarity matrix. Within each heterogeneous cluster, all clients except the cluster head perform dataset distillation and transmit the distilled data to the corresponding cluster head. The server and cluster heads then collaboratively train the model.
@@ III-C Problem Statement
In this paper, we consider a typical scenario of distributed learning with non-IID data, i.e., label distribution skew [li2022federated]. Our focus lies on addressing the issue of quantity-based label imbalance, where each client owns data with a fixed number of label classes. Supposing $N_{c}$ denotes the total number of label classes, each client owns data with $C$ different labels. Data labels among different clients could overlap. If $C=N_{c}$, the data labels among all clients are the same, which can be regarded as quantity-based label balance. If $C<N_{c}$, the data labels among all clients are imbalanced. If $C=1$, it represents the worst-case scenario of quantity-based label imbalance, where the local dataset on each client contains only a single label class. The quantity-based label imbalance has been verified to greatly decrease the performance of model training [li2020federated, geyer2017differentially].
The assumption of a fixed number of label classes per client is adopted primarily to provide a controlled and reproducible experimental setting, which is widely used in FL research for systematic evaluation of non-IID effects. In practical FL systems, however, clients may possess data from different numbers of label classes due to diverse user behaviors and data collection patterns. The proposed framework does not rely on clients having identical numbers of label classes. Instead, it leverages similarity in label knowledge inferred from soft labels to group clients and construct heterogeneous clusters, allowing it to naturally accommodate more general and realistic scenarios with arbitrary and uneven label distributions across clients.
@@ IV The Proposed HFLDD
In this section, we will first introduce the key idea of the proposed framework HFLDD. Then, we will introduce the details of each part.
@@ IV-A Overview
    Label Knowledge Collection (Step 1): Each client independently trains a local model using its private non-IID dataset. These locally trained models are then used to infer soft labels on a globally available auxiliary dataset. The resulting soft labels are uploaded to the server. This stage allows the server to gather label distribution information without accessing the raw private data, thus preserving client privacy.
@@
    IID Dataset Generation (Step 4): Within each heterogeneous cluster, all clients, except the cluster head, perform dataset distillation to generate distilled data that approximates their local data distribution. These distilled data are sent to the cluster head. The head aggregates all received distilled data, including its own data, to form an approximately IID hybrid dataset. This step helps mitigate statistical heterogeneity before training begins.
@@
    Model Training (Step 5): The server distributes the current global model to all cluster heads. Each head updates the model using the hybrid dataset and sends the updated model back to the server. The server aggregates the updated models using standard FL techniques such as FedAvg. This process is repeated for multiple rounds until model convergence.
Figure 2: An overview of heterogeneous client clustering. Each circle represents a client, with different colors representing different data labels. Initially, the server collects the soft labels uploaded by clients and calculates the similarity matrix. The server groups the clients into homogeneous clusters through K-Means. Then, heterogeneous clusters are constructed by randomly sampling clients from different homogeneous clusters, ensuring label diversity within each heterogeneous cluster.
@@ IV-B Details
The details of the implementation of each part in HFLDD are introduced in this subsection. Table II summarizes the important notations.
@@ IV-B1 Label knowledge collection
Each client first trains a local model on their own dataset $\mathcal{D}_{i}$. Since the trained local model is driven by the client dataset, it learns relevant knowledge about the local data. Subsequently, we introduce a finite and globally available dataset $\mathcal{D}_{g}=\{\left(X_{g},y_{g}\right)\}$  [^1] The global dataset $\mathcal{D}_{g}$ used for soft labels generation can be a small, publicly available dataset related to the same task but it is not necessarily representative of the client data., where $X_{g}\in\mathbb{R}^{|\mathcal{D}_{g}|\times d}$, $y_{g}\in\mathbb{R}^{|\mathcal{D}_{g}|\times N_{c}}$ and $|\mathcal{D}_{g}|$ is the number of samples in $\mathcal{D}_{g}$. The global dataset has a distribution different from the datasets of the clients and is used to generate soft labels for each client. Specifically, each client uses a globally available dataset as input for its locally pre-trained model to generate a two-dimensional tensor $S_{i}\in\mathbb{R}^{|\mathcal{D}_{g}|\times N_{c}},\forall i=1,\cdots,N$. This tensor represents the soft labels of the global dataset based on the local model of the client $i$. Let ${S}_{i,k}^{c}$ denote the probability that sample $k$ belongs to class $c$. It has been verified that the above soft labels can reflect the knowledge of the label of the local dataset [abebe2023optimizing, abebe2024improving]. Therefore, after receiving the soft labels from the clients, the server obtains the label knowledge of each client without knowing the original dataset, which prevents data privacy leakage.
@@ IV-B2 Heterogeneous Client Clustering
Based on the similarity matrix, the server employs the K-Means algorithm to cluster all the clients, by treating each row of the similarity matrix as a data point. This process groups clients with similar label knowledge into homogeneous clusters $\mathcal{H}_{o}=\left\{\mathcal{H}_{o,1},\ldots,\mathcal{H}_{o,K}\right\}$. While homogeneous clusters reveal similarity, our goal is to construct training groups with complementary label distributions. Heterogeneous clusters $\mathcal{H}_{e}=\left\{\mathcal{H}_{e,1},\ldots,\mathcal{H}_{e,H}\right\}$ are then formed by iteratively sampling across homogeneous groups to approximate inter-cluster label balance. Fig. 2 illustrates the overall process of heterogeneous client clustering. In this figure, each client’s local data contains only one label, which makes the data among different clients are severe non-IID. After clustering, each homogeneous cluster contains only one label, while each heterogeneous cluster contains labels from all the clients. The details of the sampling process are shown in Algorithm 1. The value of $K$ controls the granularity of homogeneous grouping based on label-knowledge similarity. Its role is to ensure that clients with comparable statistical characteristics are first organized into representative similarity groups, which then serve as a structured pool for constructing heterogeneous clusters. Notably, some homogeneous clusters may contain more clients than others. After smaller clusters are exhausted, the algorithm continues to sample from the remaining larger clusters. In the case where only one homogeneous cluster remains, its remaining clients are sequentially assigned to individual heterogeneous clusters. This will result in size imbalance, However, it can be mitigated by weighted aggregation and does not significantly impact convergence or model performance.
@@ IV-B3 IID dataset generation
In each heterogeneous cluster, the cluster head acts as an agent for this cluster, conducts local model training, and communicates with the server. Each cluster member, except the head, sends a distilled data of much smaller size to the cluster head. In this work, clients employ the KIP method to obtain distilled data, as introduced in Section III-A. For the $h$-th heterogeneous cluster, when the cluster head aggregates the distilled data from the entire cluster, it obtains a hybrid dataset $\mathcal{D}_{h}$, which consists of its own local data and the distilled data received from its cluster members. We transform the problem into a learning process between the server and a number of cluster heads with approximately IID data.
@@ IV-B4 Model Training
As we mentioned above, after transforming our problem into distributed learning with IID data, we can apply the classical FL method, such as FedAvg[mcmahan2017communication], for model training. In the communication round $t$, the server broadcasts the global model parameters $\omega^{t}$ to all heads of the connected cluster. Subsequently, the cluster head $h$ updates its local parameters to $\omega_{h}^{t}\,\,=\,\,\omega^{t}$, and performs $E_{2}$ local updates in $\mathcal{D}_{h}$. At the $(\tau+1)$-th epoch of local update, the local model parameters of cluster head $h$ are updated as
$\omega_{h}^{t,\tau+1}=\omega_{h}^{t,\tau}-\eta\nabla F_{h}\left(\omega_{h}^{t,\tau},\xi_{h}^{t,\tau}\right),$ (5)
where $h=1,\cdots,H$, $\tau=0,\cdots,E_{2}-1$, $\xi_{h}^{t,\tau}$ denotes a randomly selected batch from $\mathcal{D}_{h}$, $\eta$ denotes the learning rate and $\nabla F_{h}\left(\cdot\right)$ denotes the gradient of the local loss function. Finally, the server aggregates the updated local models $\omega_{1}^{t,E_{2}},...,\omega_{H}^{t,E_{2}}$ from all the cluster heads to generate a new global model.
The whole process of HFLDD is illustrated in Algorithm 2.
@@
10:for $\text{client }i=1$ to NN do
11:  for $\text{client }j=1$ to NN do
@@
[…]
@@ V-A Convergence Analysis
In HFLDD, the global training process is organized around a set of cluster heads. Each cluster head maintains a hybrid dataset and performs model training independently. The convergence behavior of HFLDD can be analyzed by extending the standard convergence theory of FedAvg with non-IID data. HFLDD maintains the following foundational assumptions commonly adopted in FL:
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Assumption 6 (Bounded distillation error): The approximation error between the empirical loss on the dataset with distillation and that without distillation at cluster $h$ is bounded:
@@
Under Assumptions 1–6, if the learning rate follows a diminishing schedule $\eta_{t}=\frac{2}{\mu(\tau+t)}$, with $\tau=\max\left\{8\frac{L}{\mu},\ E_{2}\right\}-1$, and all cluster heads participate in each communication round, then after $T$ rounds of communication, the global model $\omega^{T}$ satisfies:
[…]
@@
In HFLDD, only the cluster heads participant in global model training. It can be regarded as standard FedAvg on a cluster-level hybrid dataset $\mathcal{D}_{h}$. Based on the convergence analysis of FedAvg under non-IID conditions, as shown in [li2019convergence], the convergence bound of HFLDD satisfies:
[…]
@@ V-B Communication Cost
Now, we analyze the total communication cost of HFLDD. Considering that the initialization of the server model can be sent as a random seed [zhou2020distilled], we can ignore the communication cost of sending random seeds. Thus, we only consider the resources consumed for uploading soft labels to the server at the stage of label knowledge collection, transmitting the distilled data within each cluster at the stage of IID dataset generation, and communication between the server and cluster heads at the stage of model training.
[…]
$\mathcal{C}_{\text{HFLDD}}$ $=N\cdot|\mathcal{D}_{g}|\cdot N_{c}\cdot B_{1}+\sum_{h=1}^{H}{\sum_{\begin{subarray}{c}i\in\mathcal{H}_{e,h}\\ i\neq a_{h}\end{subarray}}{\left|\tilde{\mathcal{D}_{i}}\right|}}\cdot B_{2}$